1. A computer software retailer used a markup rate of 40%.Find the selling price of a computer game that cost the retailer P 1,250.00?



2. A golf shop pays its wholesaler P 2,000.00 for a certain club, and then sells it to a golfer for P 3, 750.00. what is the markup rate?



3. A shoe store uses a 40% markup on cost of a pair of shoes that sells for P 3,150.00?



4. A item originally priced at P 2,750.00 is marked 25% off. What is the sale price?



5. An item that regularly sells for P 3,000.00 is marked down to P2,650.00. What is the markdown rate?



6. Assume that net income equals P 10, 000 on net sales of P 100, 00. What is the profit margin?






1
Expert's answer
2021-09-13T17:11:02-0400

1.

"Markup=\%\space of \space cost \space price\\\frac{40}{100}\times1250=500\\selling \space price=500+1250\\=P1750"


2.

"cost=2000\\selling\space price =3750\\mark\space rate=\frac{sell\space price-cost\space price}{cost\space price}"


"\frac{3750-2000}{2000}\times100"


"=87.5\%"


3.

"markup=40\%\\selling\space price=3150\\\frac{40}{100}\times3150=1260"


"selling\space price=cost+markup\\=1260\times3150\\=P4,410"


4.

"markup=25\%\\cost\space price=2750\\\frac{25}{100}\times2750=6875\\sellin\space price=cost\space price +markup\\=2750+6875\\=P9,675"


5.

"cost\space price=2750\\sell\space price=3000\\marked\space down\space value=set\space price-cost\space price\\=3000-2650=350\\marked\space down\space rate=\frac{marked\space down\space value}{sell\space price}\\=\frac{350}{3000}\times100=11.67\%"


6.

"income\space10000\\net\space sales=100\\profit\space margin=\frac{income-sales}{income}\\=0.99"